OpenStudy (anonymous):

Find f'(x), f(x) = x^(1/2) (x^2 + x -4)

OpenStudy (anonymous):

\[f'(x) = (2x + 1) \cdot x^{1 \over 2} + ({1 \over 2} x ^ {-{1 \over 2}}) \cdot (x^2 + x -4)\], correct?

OpenStudy (anonymous):

f'(x)=(5/2X^3/2)+(3/2X^1/2)-(2X^-1/2)

OpenStudy (anonymous):

can you give me more details, pls?

OpenStudy (anonymous):

(fg)'=f'g+fg'

OpenStudy (anonymous):

but how you got (5/2x^3/2) ?